OpenStudy (anonymous):

Which of the following functions has the largest value when x = 3? c(x) = 3x2 + 5x + 22 j(x) = 12x a(x) = 9x All the functions are equal at x = 3. c(x) j(x) a(x)

OpenStudy (anonymous):

I think thats it is c(x)

OpenStudy (anonymous):

@Michele_Laino

OpenStudy (michele_laino):

we have to substitute x=3 into all of your equations, namely: c(3)=3*9+5*3+22=64 j(3)=12*3=36 a(3)=9*3=27 now, the largest value is 64, so you are right!

OpenStudy (anonymous):

Thank you!!!!!!!!!!

OpenStudy (michele_laino):

:)